1) x squared+5x+6=o
2) x squared-4x+3=0

Found 2 solutions by checkley71, meg1650:
Answer by checkley71(8403)

X^2+5X+6=0
(X+2)(X+3)=0
X+2=0
X=-2 SOLUTION
X+3=0
X=-3 SOLUTION
---------------------------------------
X^2-4X+3=0
(X-3)(X-1)=0
X-3=0
X=3 SOLUTION
X-1=0
X=1 SOLUTION

Answer by meg1650(2)

1) x= 2, x=3
2)x=1, x=3